Shift summary for day shift, 8/Nov/99

AG (ag@fnal.gov)
Mon, 8 Nov 1999 16:34:46 -0500 (EST)

Shift Summary for day shift, 8/Nov/99

On Shift: HW, AG

Hours of beam available at all intensities: 5.0
Hours of beam available at nominal intensity: 4.5
Hours of E799 data (special runs not included): 3.2
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:

o problems during run initialization, cleaning ktev1 tape drives.
o DC4 pre-amp power supply trip.
o Scheduled accelerator shutdown.
o TeV radiation trip.

========================================================================

Runs ended this shift:
Run 14927 - 3.3M - BEAM_799

Tapes brought to FCC this shift:
(including tapes for on-line split)

UP{A,B,C}1{09,10,11} and UPS037.

========================================================================

Accesses during shift:

a bunch of them, to fix the problems described below.

========================================================================

Overview of shift:

Rather poor E799 data-taking. Had to clean tapes on ktev1 at the beginning
of the shift, then DC4 pre-amp LV tripped at around 9:20am and most of the
triggers died/screwed up. Bunch of people went on access to fix a few
problems during the scheduled two hours accelerator shutdown (see details
below).

========================================================================

Description of Shift:

run spill time comment

------------------------------------------------------------------------

14926_______0700____Day shift begins. We are in the middle of doing dart
_______________________fresh start after run aborted earlier this
_______________________morning. Initialization takes unusually long time.
_______________________After consultation with Ed we decided to replace
_______________________tapes, clean ktev1 tape drives and do fresh start.

______________0731____Run 14927 (beam_799) started on fresh tapes
_______________________UP{A,B,C}11{5,6,7}, UPS039. (Note that tapes
_______________________UP{A,B,C}11{2,3,4} have only a muon run and
_______________________~15 minutes of beam_799 data).

______________0740____SashaL leaves.

_______49___0840_____We notice the 2Hz beats in the beam structure and
_______________________call MCR, they will work on it.

_______59___0854_____Beam is gone. LRF 4 modulator trips, OPs en-route...

_______62___0859_____Beam is back.

_______78___0921_____DC4 LV (-5V) tripped. Investigating...

_______80___0924_____There are no hyperon triggers.

____________0945______DC4 readout is dead, so for the run 14927 data is no
_______________________good after spill #78. Stopped this run. Page BobT.

____________0950______BobT walks in. There is two hours downtime scheduled
_______________________for around 1030am - pretty convenient!

____________1030______Accelerator shut down. Access in progress for
_______________________repairs and investigation.
_______________________BobT, Hogan, SashaG, JasonH, Ashkan go downstairs.

____________1120______BobT, JasonH adjusted HA offsets (see KteV log XXII,
______________________p.189).

____________1200______HN,AA,AG checked out DC4 post-amp supply. The -5V
_______________________"b" supply was dead. We checked all fuses,
_______________________connections, power etc. Turned it back on - it
_______________________looks OK now?! We decide to leave it alone for now,
_______________________but we located (with help of RickT) spare DC
_______________________post-amp in the loading dock. So if that power
_______________________supply trips again, it will be replaced.

____________1300_____RJT and AA replaced LRS 428 in HA. Also DC offsets
______________________adjusted on all channels in the linear fan in chain
______________________to ~-1mV. Details are in Muon System Log II, p.27.

____________1332_____Beam is back. Stop COMICS run.

14929______1339_____Start beam_799 run. And soon realize that NM3S is off.
______________________Called MCR. Data until spill #8 is junk.

____________1348_____Taking data. SEM ~ 8.6E12.

_______42___1435____Beam is gone. TeV radiation trip.


========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift